Stocks A and B have the following historical returns: Stock Bs Returns, le (13.40%) Year Stock As Returns, A 2014 (21.50%)Please help with part C, D, and E. Thank you!
Answer #1

c. standard deviation = √(sum of (( actual return-mean return)2 /(n-1)))

so here,

Actual return

Mean return

s= (Given return - mean)^2

year

A

B

A

B

A

B

2014

-21.50

-13.40

10.95

10.95

1,053.00

592.92

2015

39.75

26.30

10.95

10.95

829.44

235.62

2016

11.25

25.80

10.95

10.95

0.09

220.52

2017

-1.00

-13.30

10.95

10.95

142.80

588.06

2018

26.25

29.35

10.95

10.95

234.09

338.56

Total

2,259.43

1,975.69

So for 2014, for A, S= (-21.50-10.95)^2

And for B, S= (-13.40-10.95)^2

And so on

Now,

Std. dev= (s/(n-1))1/2

For A= (2259.43/4)1/2

= 564.861/2

= 23.77%

For B= (1975.69/4)1/2

= 493.921/2

= 22.22%

Now, for the portfolio:

Portfolio return

mean return

s= (Given return - mean)^2

year

A

2014

-17.45

10.95

806.56

2015

33.03

10.95

487.31

2016

18.53

10.95

57.38

2017

-7.15

10.95

327.61

2018

27.80

10.95

283.92

Total

1,962.78

Std. dev= (s/(n-1))1/2

= (1962.78/4)^1/2

= 490.69^1/2

= 22.15%

d. Coefficient of variation:

Coefficient of variation (COV)= standard deviation/ expected return rate or mean

for A:

SD= 23.77

Mean= 10.95

COV= 23.77/10.95

= 2.17

for B:

SD= 22.22

Expected return = 10.95

COV= 22.22/10.95

=2.03

For the portfolio:

SD= 22.15

Mean= 10.95

COV= 22.15/10.95

= 2.02

e. as a risk aversive investor, we will look into the stock that has the minimum deviation from the expected return , that is we will go for the portfolio of 50% A and 50% B.